# PET: Preference Evolution Tracking with LLM-Generated Explainable Distribution

### v2

@@ Appendix D Proofs
By 1, the scores $S$ from PET preserve the order of the true latent attractiveness scores $\bm{q}$. Since the softmax function is monotonic, the inferred probability distribution $\hat{\theta}$ also preserves this order. Thus, sorting by PET’s inferred probabilities produces a ranking, $\pi_{\text{PET}}$, that is identical to the optimal ranking, $\pi^{*}$, which sorts by the true latent scores $\bm{q}$. For a given user and within the time window of $\{t+1:T\}$, we let the vector of attractiveness score $\bm{q}$ be the true latent relevance score of a user for the cluster space. And we denote the sorted (in descending order) vector of $\bm{q}$ as $\tilde{\bm{q}}$ with $\tilde{q}_{1}\geq\cdots\geq\tilde{q}_{K}$. By 1, the retried logit score vector preserves the same order $S$ as $\bm{q}$. Thus, as the softmax function of $S$ preserves the same order of $S$, then by the procedure of PET – sorting the inferred probability distribution in the descending order – the outputted rank by the PET’s procedure has the exactly the same order as the $\tilde{\bm{q}}$.
On the other hand, let’s visit the definition of three ranking metrics. For any given ranking/permutation of cluster set $C$, denoted as $\pi$ (with the associated score vector $q_{\pi}$), we define
  • Recall@k: $\{\text{relevant items in top}-k\}/K$.
@@
  • NDCG@k: $\text{DCG}(\pi)@k/\text{IDCG}@k$, where $\text{DCG}(\pi)@k:=\sum_{r+1}^{K}\frac{q_{\pi}(r)}{\log_{2}(r+1)}$ and $\text{IDCG}@k$ is the best possible $\text{DCG}(\pi)@k$ over all possible permutation $\pi$.
It is trivial to see that, all these three ranking metrics are maximized when the permutation $\pi$ has the same order as $\tilde{\bm{q}}$. Since the outputted rank by PET has the same order as $\tilde{\bm{q}}$, then PET achieves the optimal ranking scores with respect to these three ranking methods, which gives equal or better rank scores than the direct generation method. ∎
@@ Appendix E Detailed Experimental Design
E.1 Yelp Dataset
On the Yelp dataset, we evaluate PET in a high-dimensional setting with 1,311 fine-grained categories in the Yelp dataset, designed to test the scalability of our approach. In this regime, Logit Probing becomes prohibitively expensive due to the large output space. To address this, we adopt Hierarchical Probing (Algorithm 3) that decomposes the task into two levels: a coarse-grained cluster identification followed by fine-grained ranking within the selected scope.
We construct a hierarchical mapping that maps 1,311 categories into 26 high-level L1 clusters (e.g., “Food & Restaurants”, “Health & Medical”). For a high-traffic category like “Food & Restaurants,” we then define a corresponding L2 space with 19 sub-clusters (e.g., “Mexican,” “Vegan”) under L1’s “Food & Restaurants”. In inference, PET first identifies the user’s broad L1 interests and then “zooms in” to perform a nuanced L2 ranking within the top-predicted L1 category. While operators may freely select which high-level categories to probe based on application needs, this experiment offers one representative setup on the Yelp dataset.
@@ E.1 Yelp Dataset
L1 Categories: Active Life & Fitness, Active Life, Sports & Recreation, Arts, Entertainment & Events, Automotive, Beauty & Spas, Cannabis Services, Community & Government, Education, Event Planning & Services, Farms & Ranches, Food & Restaurants, Health & Medical, Home & Public Services, Home Maintenance, Hotels & Travel, Internet & Communications, Local & Public Services, Nightlife & Bars, Personal Services, Pets, Professional & Financial Services, Real Estate, Religious & Community, Shopping & Retail, Specialty Shops, Specialty Vehicles L2 Categories: Bicycles, Bike Shop, Active Life, Aerial Fitness, Airsoft, Amateur Sports Teams, Archery, Badminton, Barre Classes, Baseball Fields, Basketball Courts, Beach Bars, Beaches, Bicycle Paths, Bike Parking, Bike Sharing, Bike tours, Bikes, Bocce Ball, Boot Camps, Boxing, Brazilian Jiu-jitsu, …
E.2 Real-World Short-Video Dataset
@@ Appendix F Additional Experimental Results
F.1 Evaluation of PET on Movielens
Inference procedure of Direct Generation. Unlike PET, which infers a full probability distribution over all 19 clusters, Direct Generation is prompted separately for each top-k values.

### v3

@@ D.1 Preliminaries: Idealized Bayes-optimality
Let the true latent preference distribution be $\theta=\mathrm{softmax}(\boldsymbol{q})$ and the predicted distribution be $\hat{\theta}=\mathrm{softmax}(\boldsymbol{S})$. The expected population cross-entropy risk is defined as:
$\mathcal{R}=\mathbb{E}_{H}\left[-\sum_{k=1}^{K}\theta_{k}\log\hat{\theta}_{k}\right].$ (3)
By adding and subtracting the entropy of the true distribution $H(\theta)=-\sum\theta_{k}\log\theta_{k}$, we can rewrite the risk as:
@@
where $D_{\mathrm{KL}}$ is the Kullback-Leibler divergence. A fundamental property of proper scoring rules (Gibbs’ inequality) states that $D_{\mathrm{KL}}(\theta\|\hat{\theta})\geq 0$, with equality holding if and only if $\hat{\theta}=\theta$ almost everywhere. Therefore, the global minimum is achieved uniquely when $\mathrm{softmax}(\boldsymbol{S})=\mathrm{softmax}(\boldsymbol{q})$.
The softmax function is invariant to constant shifts, meaning $\mathrm{softmax}(\boldsymbol{x})=\mathrm{softmax}(\boldsymbol{y})$ implies $\boldsymbol{x}=\boldsymbol{y}+c\boldsymbol{1}$ for some scalar $c\in\mathbb{R}$. Consequently, at the global minimum, $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$.
Finally, since translation by a scalar $c$ does not alter the relative ordering of elements, for any pair of clusters $i,j$:
$S_{i}>S_{j}\iff(q_{i}+c)>(q_{j}+c)\iff q_{i}>q_{j}.$ (5)
This confirms that in the idealized limit, the logits $\boldsymbol{S}$ are perfectly isotonic with the latent utility $\boldsymbol{q}$. ∎
In the language-modeling context, this corresponds to the standard interpretation of cross-entropy training as maximum-likelihood estimation of the next-token distribution and an upper bound on the true entropy rate (Braverman et al., 2020). Our analysis applies this Bayes-optimal perspective to the cluster-distribution predictor used by SPECTRA.
D.2 Robustness: The $\varepsilon$-Approximate Regime
Proposition proposition˜1 characterizes an ideal Bayes-optimal limit where $\boldsymbol{S}$ is exactly isotonic with $\boldsymbol{q}$. In practice, models are only approximately calibrated (Blasiok et al., 2023; Braverman et al., 2020). We capture this via a probabilistic pairwise condition.
Assumption 1 ($\varepsilon$-Approximate Pairwise Isotonicity).
For a given user and time window, let $\boldsymbol{q}\in\mathbb{R}^{K}$ be latent scores and $\boldsymbol{S}\in\mathbb{R}^{K}$ the logits produced by SPECTRA. We say $\boldsymbol{S}$ is $\varepsilon$-approximately pairwise isotonic w.r.t. $\boldsymbol{q}$ if, for all $i,j$ with $q_{i}>q_{j}$,
$\mathbb{P}\Big(S_{i}<S_{j}\Big)\;\leq\;\varepsilon,$ (6)
where the probability is over randomness in training, histories, and probing. That is, each strictly preferred pair is misordered with probability at most $\varepsilon$.
We measure ranking quality by the number of misordered pairs relative to the Bayes-optimal ranking.
Definition 1 (Pairwise ranking regret).
Let $\pi^{*}$ be the permutation that sorts clusters by decreasing $q_{i}$, and let $\pi$ be any permutation of $\{1,\dots,K\}$. Define
$\mathcal{R}(\pi;\boldsymbol{q})=\sum_{\begin{subarray}{c}i,j\in[K]\\ q_{i}>q_{j}\end{subarray}}\mathbb{I}\big(\pi^{-1}(i)>\pi^{-1}(j)\big),$ (7)
where $\pi^{-1}(i)$ denotes the position (rank) of cluster $i$ under the permutation $\pi$ (smaller is better). Thus, $\mathcal{R}(\pi;\boldsymbol{q})$ counts the number of strictly preferred pairs $(i,j)$ that are reversed under $\pi$.
Let $\pi_{\mathrm{SPECTRA}}$ be the permutation that sorts $\boldsymbol{S}$ in descending order. The following bound is immediate.
Theorem 2 (Approximate optimality of SPECTRA).
Under Assumption 1, the expected pairwise regret of $\pi_{\mathrm{SPECTRA}}$ satisfies
$\mathbb{E}\big[\mathcal{R}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})\big]\;\leq\;\binom{K}{2}\,\varepsilon=O(K^{2}\varepsilon),$ (8)
where the expectation is over the randomness in $\boldsymbol{S}$.
Proof.
For $q_{i}>q_{j}$, define $E_{ij}=\{\pi_{\mathrm{SPECTRA}}^{-1}(i)>\pi_{\mathrm{SPECTRA}}^{-1}(j)\}=\{S_{i}<S_{j}\}.$ By Assumption 1, $\mathbb{P}(E_{ij})\leq\varepsilon$. By Definition 1, $\mathcal{R}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})=\sum_{q_{i}>q_{j}}\mathbb{I}(E_{ij}).$ Taking expectations and using linearity,
$\mathbb{E}[\mathcal{R}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})]=\sum_{q_{i}>q_{j}}\mathbb{P}(E_{ij})\leq\sum_{q_{i}>q_{j}}\varepsilon\leq\binom{K}{2}\varepsilon.$ (9)
Thus, if SPECTRA is “mostly isotonic” in the sense of Assumption 1 (small $\varepsilon$), its ranking is close to Bayes-optimal: only $O(K^{2}\varepsilon)$ pairs are misordered in expectation. In the limit $\varepsilon\to 0$, we recover the exact isotonicity of Proposition 1.
D.3 From cross-entropy to the $\varepsilon$ regime
The previous result treats $\varepsilon$ as a behavioral property of the SPECTRA head. We now connect $\varepsilon$ to the training objective $\mathcal{L}_{\mathrm{CE}}(\phi)$.
Let $\mathcal{L}_{\mathrm{CE}}^{\star}=\mathcal{L}_{\mathrm{CE}}(\theta_{t+1:T})$ denote the Bayes-optimal risk, achieved when $\hat{\theta}_{t+1:T}=\theta_{t+1:T}$, and define the excess cross-entropy risk
$\Delta_{\mathrm{CE}}$ $=\mathcal{L}_{\mathrm{CE}}(\phi)-\mathcal{L}_{\mathrm{CE}}^{\star}$ (10)
$=\mathbb{E}_{H_{1:t}}\big[\mathrm{KL}\big(\theta_{t+1:T}(H_{1:t})\,\|\,\hat{\theta}_{t+1:T}(H_{1:t})\big)\big].$ (11)
By Pinsker’s inequality and Jensen’s inequality, we obtain:
Lemma 1 (L1 error from excess cross-entropy).
The expected $L_{1}$ distance between the latent and predicted preference distributions is bounded by
$\mathbb{E}_{H_{1:t}}\big[\|\theta_{t+1:T}(H_{1:t})-\hat{\theta}_{t+1:T}(H_{1:t})\|_{1}\big]\;\leq\;$
$\quad\sqrt{2\,\Delta_{\mathrm{CE}}}.$ (12)
Proof.
Recall that the excess cross-entropy risk is exactly the expected Kullback-Leibler divergence between the true distribution $\theta$ and the predicted distribution $\hat{\theta}_{t+1:T}$:
[…]
For any specific history $H_{1:t}$, Pinsker’s inequality bounds the $L_{1}$ distance by the KL divergence:
$\|\theta-\hat{\theta}\|_{1}\;\leq\;\sqrt{2\,D_{\mathrm{KL}}(\theta\|\hat{\theta})}$. (14)
Taking the expectation over histories $H_{1:t}$ on both sides:
$\mathbb{E}\left[\|\theta-\hat{\theta}\|_{1}\right]\;\leq\;\mathbb{E}\left[\sqrt{2\,D_{\mathrm{KL}}(\theta\|\hat{\theta})}\right$]. (15)
Since the square root function $f(x)=\sqrt{x}$ is concave, Jensen’s inequality implies that $\mathbb{E}[\sqrt{X}]\leq\sqrt{\mathbb{E}[X]}$. Applying this to the right-hand side:
$\mathbb{E}\left[\sqrt{2\,D_{\mathrm{KL}}(\theta\|\hat{\theta})}\right$]\; $\leq\;\sqrt{2\,\mathbb{E}\left[D_{\mathrm{KL}}(\theta\|\hat{\theta})\right]}\;$ (16)
$=\;\sqrt{2\,\Delta_{\mathrm{CE}}}.$ (17)
Combining these steps yields the lemma. ∎
As $\Delta_{\mathrm{CE}}\to 0$ (e.g., with more data, capacity, and optimization), the predicted distributions $\hat{\theta}_{t+1:T}$ converge to $\theta_{t+1:T}$ in expected $L_{1}$ distance. Intuitively, if $\hat{\theta}_{t+1:T}$ is close to $\theta_{t+1:T}$, then it is unlikely to significantly distort the ordering between clusters with a clear preference gap. Formalizing $\varepsilon$ as an explicit function of $\Delta_{\mathrm{CE}}$ for all pairs would require additional global margin assumptions on $\theta_{t+1:T}$, which are often unrealistic in long-tail regimes. Instead, we interpret $\varepsilon$ as capturing residual misorderings after training: Lemma˜1 shows that cross-entropy minimization drives $\hat{\theta}_{t+1:T}$ toward $\theta_{t+1:T}$, thereby pushing the model into a small-$\varepsilon$ regime for most histories and clearly preferred pairs.
D.3  SPECTRA vs. Decoding-Based Direct Generation
We now compare SPECTRA to decoding-based direct generation (DG), which uses the same logits $\boldsymbol{S}$ but produces a ranked list via an autoregressive decoding algorithm (e.g., greedy, beam, sampling).
Let $\pi^{*}\in\Pi_{K}$ be the permutation that sorts clusters by decreasing $\boldsymbol{q}$, and let $\pi_{\mathrm{SPECTRA}}$ be the permutation that sorts $\boldsymbol{S}$. We model any decoding-based ranking as follows.
Definition 2 (Decoding-based ranking).
A decoding-based ranking procedure is a (possibly randomized) algorithm that maps logits $\boldsymbol{S}\in\mathbb{R}^{K}$ to a permutation of clusters. We denote by
$\pi_{\mathrm{DG}}(\boldsymbol{S})\in\Pi_{K}$
the (random) permutation produced by such a decoder when run on $\boldsymbol{S}$. The randomness here comes from any sampling or tie-breaking used by the decoding algorithm (e.g., top-$p$ sampling, temperature, stochastic beam search). We write $\mathbb{E}_{\mathrm{DG}}[\cdot]$ for expectation with respect to this decoder randomness, conditioned on the logits $\boldsymbol{S}$.
We measure quality using a generic order-aware loss.
Assumption 2 (Order-aware ranking loss).
Let $\mathcal{L}:\Pi_{K}\times\mathbb{R}^{K}\to\mathbb{R}$ be a loss such that, for any fixed $\boldsymbol{q}$:
  1. 1.
    $\mathcal{L}(\pi^{*};\boldsymbol{q})\leq\mathcal{L}(\pi;\boldsymbol{q})$ for all $\pi\in\Pi_{K}$;
  2. 2.
    for almost all $\boldsymbol{q}$ (w.r.t. any continuous distribution), the minimizer is unique: $\mathcal{L}(\pi^{*};\boldsymbol{q})<\mathcal{L}(\pi;\boldsymbol{q})$ for all $\pi\neq\pi^{*}$.
In the Bayes-optimal regime of Proposition 1, $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$, SPECTRA recovers $\pi^{*}$ by simple sorting. We show that no decoding procedure can do better on top of these logits.
Theorem 3 (Dominance of SPECTRA over decoding-based generation).
Fix $\boldsymbol{q}\in\mathbb{R}^{K}$ and assume $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$ for some $c$. Let $\pi_{\mathrm{SPECTRA}}=\mathrm{argsort}\boldsymbol{S}$ and let $\pi_{\mathrm{DG}}(\boldsymbol{S})\in\Pi_{K}$ be the (possibly random) permutation produced by any decoding-based procedure when run on logits $\boldsymbol{S}$. Under Assumption 2,
$\mathbb{E}_{\mathrm{DG}}\big[\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})\big]\;\geq\;\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q}),$ (18)
where the expectation is over the decoder’s internal randomness. Moreover, if $\mathcal{L}(\cdot;\boldsymbol{q})$ has a unique minimizer $\pi^{*}$ and the decoder outputs any $\pi\neq\pi^{*}$ with nonzero probability, then the inequality is strict.
Proof.
Since $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$, sorting $\boldsymbol{S}$ is equivalent to sorting $\boldsymbol{q}$: $\pi_{\mathrm{SPECTRA}}=\mathrm{argsort}\boldsymbol{S}=\mathrm{argsort}\boldsymbol{q}=\pi^{*}.$ By Assumption 2, $L(\pi^{*};\boldsymbol{q})\leq L(\pi;\boldsymbol{q})$ for all $\pi\in\Pi_{K}$.
For any fixed realization of the decoder’s randomness, the decoder outputs some permutation $\pi_{\mathrm{DG}}(\boldsymbol{S})\in\Pi_{K}$, and thus
$\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})\;\geq\;\mathcal{L}(\pi^{*};\boldsymbol{q})=\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q}).$ (19)
Taking expectation over the decoder’s randomness gives
$\mathbb{E}_{\mathrm{DG}}\big[\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})\big]\;\geq\;\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q}),$ (20)
since $\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})$ is deterministic.
If $\mathcal{L}(\cdot;\boldsymbol{q})$ has a unique minimizer $\pi^{*}$, then whenever $\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})=\mathcal{L}(\pi^{*};\boldsymbol{q})$ we must have $\pi_{\mathrm{DG}}(\boldsymbol{S})=\pi^{*}$. Thus equality in expectation can hold only if the decoder returns $\pi^{*}$ almost surely. If instead the decoder outputs some $\pi\neq\pi^{*}$ with nonzero probability, then on that event $\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})>\mathcal{L}(\pi^{*};\boldsymbol{q})$, which makes the inequality strict. ∎
Theorem 3 is an idealized statement: when logits are perfectly calibrated ($\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$), SPECTRA’s simple logit-sorting is Bayes-optimal, and any decoding-based mapping on top of the same logits can at best match, but not improve, the induced ranking. In practice, SPECTRA operates in the $\varepsilon$-approximate regime of Assumption 1, where its regret is controlled by Theorem 2, while decoding-based methods incur additional errors from exposure bias, truncation, and search heuristics.
Remark 1 (Risk Decomposition: SPECTRA vs. Direct Generation).
Theorem 3 characterizes an idealized regime where logits are perfectly calibrated ($\boldsymbol{S}\approx\boldsymbol{q}$). In the practical regime, Direct Generation (DG) incurs additional error sources that SPECTRA avoids due to its distributional nature. We formally conceptualize the ranking risk of DG as:
$\mathrm{Risk}(\text{DG})$ $\;\approx\;\mathrm{Risk}(\text{SPECTRA})\;+\;\Delta_{\text{freq}}$
$\quad\;+\;\Delta_{\text{trunc}}\;+\;\Delta_{\text{exp}}.$ (21)
Here, $\mathrm{Risk}(\text{SPECTRA})$ is controlled by the approximate isotonicity parameter $\varepsilon$ (Theorem 2), while the additive penalty terms represent the structural limitations identified in Section 3.3:
  1. 1.
    $\Delta_{\text{freq}}$ (Frequency Bias): Represents the divergence caused by the autoregressive model’s tendency to over-produce high-frequency tokens from pre-training (Zhao et al., 2021; Holtzman et al., 2020), which SPECTRA mitigates via closed-set normalization.
  2. 2.
    $\Delta_{\text{trunc}}$ (Search/Truncation Error): Represents the expected utility loss from heuristic search (e.g., beam search), which imposes a hard truncation on valid long-tail clusters falling outside the beam (Wiseman and Rush, 2016; Stahlberg and Byrne, 2019).
  3. 3.
    $\Delta_{\text{exp}}$ (Exposure Bias): Represents the error accumulated due to context distribution shift during autoregressive decoding (Ranzato et al., 2015).
This decomposition theoretically grounds our empirical observation that SPECTRA is significantly more robust in long-tail ranking tasks.
@@ Appendix E Detailed Experimental Design
E.1 Setups
Global Exposure Entropy. Let $\mathcal{U}$ be the set of evaluation users and $\mathcal{C}$ the cluster space (e.g., genres). For a given method and cutoff $k$, let $R_{u,k}\subseteq\mathcal{C}$ denote the set of clusters appearing in the top-$k$ list for user $u\in\mathcal{U}$. We define the global exposure count for cluster $c\in\mathcal{C}$ as
$n_{c}(k)=\sum_{u\in\mathcal{U}}\mathbb{I}\big(c\in R_{u,k}\big),$ (22)
and the corresponding normalized exposure distribution
$P_{\mathrm{exp},k}(c)=\frac{n_{c}(k)}{\sum_{c^{\prime}\in\mathcal{C}}n_{c^{\prime}}(k)}.$ (23)
The Global Exposure Entropy at cutoff $k$ is then
$\mathrm{Entropy@}k=H_{k}:=-\sum_{c\in\mathcal{C}}P_{\mathrm{exp},k}(c)\,\log_{2}P_{\mathrm{exp},k}(c).$ (24)
Higher values of $H_{k}$ indicate that exposure is distributed more equitably across clusters, while lower values indicate concentration on a small subset of clusters (mode collapse). We use presence in the top-$k$ list (unweighted) to measure exposure concentration; position-weighted variants yield similar trends.
E.2 Yelp Dataset
On the Yelp dataset, we evaluate SPECTRA in a high-dimensional setting with 1,311 fine-grained categories in the Yelp dataset, designed to test the scalability of our approach. In this regime, Logit Probing becomes prohibitively expensive due to the large output space. To address this, we adopt Hierarchical Probing (Algorithm˜3) that decomposes the task into two levels: a coarse-grained cluster identification followed by fine-grained ranking within the selected scope.
We construct a hierarchical mapping that maps 1,311 categories into 26 high-level L1 clusters (e.g., “Food & Restaurants”, “Health & Medical”). For a high-traffic category like “Food & Restaurants,” we then define a corresponding L2 space with 19 sub-clusters (e.g., “Mexican,” “Vegan”) under L1’s “Food & Restaurants”. In inference, SPECTRA first identifies the user’s broad L1 interests and then “zooms in” to perform a nuanced L2 ranking within the top-predicted L1 category. While operators may freely select which high-level categories to probe based on application needs, this experiment offers one representative setup on the Yelp dataset.
@@ E.2 Yelp Dataset
L1 Categories: Active Life & Fitness, Active Life, Sports & Recreation, Arts, Entertainment & Events, Automotive, Beauty & Spas, Cannabis Services, Community & Government, Education, Event Planning & Services, Farms & Ranches, Food & Restaurants, Health & Medical, Home & Public Services, Home Maintenance, Hotels & Travel, Internet & Communications, Local & Public Services, Nightlife & Bars, Personal Services, SPECTRAs, Professional & Financial Services, Real Estate, Religious & Community, Shopping & Retail, Specialty Shops, Specialty Vehicles L2 Categories: Bicycles, Bike Shop, Active Life, Aerial Fitness, Airsoft, Amateur Sports Teams, Archery, Badminton, Barre Classes, Baseball Fields, Basketball Courts, Beach Bars, Beaches, Bicycle Paths, Bike Parking, Bike Sharing, Bike tours, Bikes, Bocce Ball, Boot Camps, Boxing, Brazilian Jiu-jitsu, …
E.3 Real-World Short-Video Dataset
@@ Appendix F Additional Experimental Results
F.1 Evaluation of SPECTRA on Movielens
Figure 4: Comprehensive comparison of different methods on MovieLens dataset across a range of context windows (i.e., 1,3,5, and 8 sessions). Prediction windows: long-term and short-term. Metrics: NDCG@$1,5,10$ and JS-Divergence. We compare our Logit Probing and Generative Classification methods against (1) Direct Generation benchmark trained (CPT+SFT) on a Qwen3-8B base model and (2) a SOTA Qwen3-Reranker-8B ranking model. Note that for the Direct Generation baseline, the JS-divergence will not be available. Sample size: 38,434.
Qwen3-8B DeepSeek-distill-Llama-8B
Prediction Method Entropy@1 Entropy@3 Entropy@5 Entropy@10 Entropy@1 Entropy@3 Entropy@5 Entropy@10
Long A: Direct Generation 1.020 2.446 3.558 3.849 0.140 1.629 2.666 3.447
B: Logit-Probing 1.747 3.130 3.565 3.979 2.307 3.054 3.352 3.813
C: Gen. Class. 1.575 2.732 3.103 3.705 0.952 1.926 2.608 3.329
Short A: Direct Generation .810 2.163 3.581 3.767 0.050 1.635 2.713 3.476
B: Logit-Probing 2.093 3.246 3.612 3.964 2.256 2.922 3.252 3.701
C: Gen. Class. 1.605 2.696 3.211 3.681 0.073 1.972 2.350 3.433
Table 6: Fairness performance on Movielens using the CPT+SFT pipeline. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: entropy score; context window: 3. Best results for each setting are in bold.
Inference procedure of Direct Generation. Unlike SPECTRA, which infers a full probability distribution over all 19 clusters, Direct Generation is prompted separately for each top-k values.